OpenStudy (anonymous):

Find dy/dx by implicit differentiation. tan(x-y)=y/(2+x^2)

OpenStudy (anonymous):

\[\sec^2(x-y)(1-y')=\frac{(2+x^2)y'-y(2x)}{(2+x^2)^2}\] is a start

OpenStudy (anonymous):

then a raft of not very interesting algebra to solve this equation for \(y'\)

OpenStudy (anonymous):

that's the part that's tripping me up, the algebra

OpenStudy (anonymous):

how do i go about getting y' on one side

OpenStudy (anonymous):

just gotta grind it out

OpenStudy (anonymous):

distribute first

OpenStudy (anonymous):

haha don't know where to start

OpenStudy (anonymous):

\[\sec^2(x-y)(1-y')=\frac{(2+x^2)y'-y(2x)}{(2+x^2)^2}\] \[\sec^2(x-y)-y'\sec^2(x-y)=\frac{(2+x^2)y'-2xy)}{(2+x^2)^2}\]

OpenStudy (anonymous):

\[\sec^2(x-y)-y'\sec^2(x-y)=\frac{y'}{2+x^2}-\frac{2xy}{(2+x^2)^2}\]

OpenStudy (anonymous):

\[\sec^2(x-y)+\frac{2xy}{(2+x^2)^2}=\frac{y'}{2+x^2}+y'\sec^2(x-y)\]

OpenStudy (anonymous):

\[\sec^2(x-y)+\frac{2xy}{(2+x^2)^2}=y'\left(\frac{1}{2+x^2}+\sec^2(x-y)\right)\]

OpenStudy (anonymous):

thanks! this helped